OpenStudy (anonymous):

HELP! How would I find the correct simplification 8^2 x 8^-5

OpenStudy (anonymous):

\[8^{2} * 8 ^{-5}\] both have exponents, so you have to check Laws of exponents \[a ^{m} * a ^{n} = a ^{m + n}\] can you see both have the same base = 8 now lets do as the Law says: \[8^{2 - 5}\] \[= 8^{-3}\] Another Law of exponents: \[a ^{-m} = \frac{ 1 }{ a ^{m} }\] you can see that our base of 8 has a negative number so lets lets apply our law. \[\frac{ 1 }{ 8^{3} }\] or \[\frac{ 1 }{ 512 }\] hope this helps